Every month, there are 1000 independent TIE fighter flights, and each TIE fighter flight crashes with a probability of 0.0035.

(a) What is the probability that at least 2 crashes occur in the next month? You do not need to evaluate the sum you write down.

(b) The sum you wrote in Part (a) is ridiculous to evaluate. Instead, approximate the value by using a Poisson distribution. Do not leave your answer as an infinite sum here.

(c) Briefly explain why a Poisson distribution here is a good choice for approximation.
Answer #1

solution :- xn Binomial (1000) n-x 0.0035) P(x) = (♥) px qm Plx29) = 1- 3 P(x=0) +PIX=} E1-77100)(0.0035) (1–0-0035) (1009) (